# Dynamic Rank Adjustment for Accurate and Efficient Neural Network Training

### v1

@@ 2 Background
Learning Rate and Noise Scale – In previous works, learning rate is known to play a key role in determining generalization performance of machine learning models. It has been theoretically shown that the noise scale $g$ in gradient approximation is determined by the learning rate and batch size, as follows [27]: $g\approx\eta N/B$, where $\eta$ is learning rate, NN is the dataset size, and $B$ is the mini-batch size. This relation links learning rate schedules to the notion of noise scale. In early training, a large noise scale helps the model explore and shape the decision boundary [22, 21], consistent with the common practice of decaying the learning rate. In this study, we investigate the relationship between the noise scale and the effective rank of the model’s weights.
@@ 4.2 Model Rank Scheduling
Let the total training budget be $E$ epochs. We begin training in a low-rank form and increase the rank of model weights at epoch II. Then, at epoch $D$, we reduce the rank again. That is, training is performed in the full-rank space for DID-I epochs, and in the low-rank space for the remaining E(DI)E-(D-I) epochs. Under this setting, the key question is how to find good II and $D$. Here, we present two empirical insights that motivate practical principles for selecting II and $D$ to maximize rank recovery.
@@
Table 2: Layer-wise comparison of singular-value spectral ratio ($\lambda$) across different rank-adjustment schedules. The II and $D$ indicate rank inflation and deflation, respectively. ResNet20 was trained on CIFAR-10 for 150 epochs, with the learning rate reduced by a factor of 10 at epochs 100 and 130. Having full-rank epochs at the end of the high-noise regime yields the best results.
@@ 4.3 Unified Rank Adjustment Framework
Based on the two key insights discussed above, we build up a general model rank adjustment framework as shown in Figure 2. The training begins in a low-rank space. The model’s rank is adjusted twice throughout the training, increased in a high-noise regime and then decreased in a low-noise regime. In our empirical study, we observed reasonably good performance when II is set to the middle of high-noise regime and $D$ to the middle of low-noise regime. For example, given $E=150$, if learning rate decays at epoch $100$, we recommend beginning with $I=50$ and $D=125$ and then fine-tune these values to minimize training time while maintaining model accuracy. We will include an ablation study to analyze how the choices of II and $D$ affect model accuracy and computational cost in the following section.
@@ 5.1 Comparative Study
Table 4 reports the results. We compare our method with LoRA [10], the most widely used low-rank fine-tuning method, and AdaLoRA [33], a representative variant. All low-rank methods greatly reduce compute versus full fine-tuning. LoRA is the most efficient but shows a noticeable accuracy drop. AdaLoRA mitigates the drop but still does not match the full fine-tuning accuracy. By contrast, dynamic-rank training achieves accuracy comparable to that of full fine-tuning. Although the computational cost is roughly 77 times AdaLoRA’s cost but still far cheaper than full fine-tuning, making it a practical choice for NLP applications.
@@ 5.2 Ablation Study
In our empirical study, we found that dynamic-rank training performs well when $D$ and II are set to allocate a similar number of full-rank epochs to both the high-noise and low-noise regimes. We now conduct a simple ablation study to examine how the number of full-rank epochs affects model accuracy. Table 6 presents the results. SVD-based low-rank training is evaluated on the CIFAR-10 and CIFAR-100 benchmarks. For convenience, we define $\phi=(D-I)/E$ as the ratio of full-rank epochs to the total training budget. When $\phi=0.0$, it corresponds to conventional low-rank training; when $\phi=1.0$, it becomes full-rank training. On both benchmarks, accuracy starts to drop when $\phi$ goes below 0.5, particularly in the range $\phi\in[0.3,0.5]$. We therefore recommend starting with $\phi=0.5$ and adjusting downward as needed.
@@
Table 6: Ablation study on how $\phi=(D-I)/E$ affects the performance of dynamic-rank training. Low-Rank SVD is used for all experiments. When $\phi=0.0$, it becomes the conventional low-rank training. In contrast, when $\phi=1.0$, it becomes full-rank training. The II and $D$ are set to perform the same number of full-rank epochs in the high-noise and low-noise regimes.
@@ 5.3 Computational Cost Analysis
Thus, if $\rho$ is the same, $T_{DR}$ should be higher than $T_{SVD}$ and lower than $T_{F}$. For example, ResNet20 contains 272,762 trainable parameters and the reparameterized model with SVD ($\rho=0.5$) has 155,170 parameters. When $E=150$, $I=60$, and $D=135$, where $\phi=0.5$, $T_{DR}/T_{F}$ becomes $0.7844$, which is shown in Comp. column in Table 3. Our empirical study demonstrates that, by carefully tuning II and $D$, a relatively small DID-I value provides comparable accuracy to the full-rank training. Small DID-I settings in dynamic-rank training makes its training cost more similar to that of the conventional low-rank training.

### v2

@@ 2 Background
Learning Rate and Noise Scale – In previous works, learning rate is known to play a key role in determining generalization performance of machine learning models. It has been theoretically shown that the noise scale $g$ in gradient approximation is determined by the learning rate and batch size, as follows [27]: $g\approx\eta N/B$, where $\eta$ is learning rate, $N$ is the dataset size, and $B$ is the mini-batch size. This relation links learning rate schedules to the notion of noise scale. In early training, a large noise scale helps the model explore and shape the decision boundary [22, 21], consistent with the common practice of decaying the learning rate. In this study, we investigate the relationship between the noise scale and the effective rank of the model’s weights.
@@ 4.2 Model Rank Scheduling
Let the total training budget be $E$ epochs. We begin training in a low-rank form and increase the rank of model weights at epoch $I$. Then, at epoch $D$, we reduce the rank again. That is, training is performed in the full-rank space for DID-I epochs, and in the low-rank space for the remaining E(DI)E-(D-I) epochs. Under this setting, the key question is how to find good $I$ and $D$. Here, we present two empirical insights that motivate practical principles for selecting $I$ and $D$ to maximize rank recovery.
@@
Table 2: Layer-wise comparison of singular-value spectral ratio ($\lambda$) across different rank-adjustment schedules. The $I$ and $D$ indicate rank inflation and deflation, respectively. ResNet20 was trained on CIFAR-10 for 150 epochs, with the learning rate reduced by a factor of 10 at epochs 100 and 130. Having full-rank epochs at the end of the high-noise regime yields the best results.
@@ 4.3 Unified Rank Adjustment Framework
Based on the two key insights discussed above, we build up a general model rank adjustment framework as shown in Figure 2. The training begins in a low-rank space. The model’s rank is adjusted twice throughout the training, increased in a high-noise regime and then decreased in a low-noise regime. In our empirical study, we observed reasonably good performance when $I$ is set to the middle of high-noise regime and $D$ to the middle of low-noise regime. For example, given $E=150$, if learning rate decays at epoch $100$, we recommend beginning with $I=50$ and $D=125$ and then fine-tune these values to minimize training time while maintaining model accuracy. We will include an ablation study to analyze how the choices of $I$ and $D$ affect model accuracy and computational cost in the following section.
@@ 5.1 Comparative Study
Table 4 reports the results. We compare our method with LoRA [10], the most widely used low-rank fine-tuning method, and AdaLoRA [33], a representative variant. All low-rank methods greatly reduce compute versus full fine-tuning. LoRA is the most efficient but shows a noticeable accuracy drop. AdaLoRA mitigates the drop but still does not match the full fine-tuning accuracy. By contrast, dynamic-rank training achieves accuracy comparable to that of full fine-tuning. Although the computational cost is roughly $7$ times AdaLoRA’s cost but still far cheaper than full fine-tuning, making it a practical choice for NLP applications.
@@ 5.2 Ablation Study
In our empirical study, we found that dynamic-rank training performs well when $D$ and $I$ are set to allocate a similar number of full-rank epochs to both the high-noise and low-noise regimes. We now conduct a simple ablation study to examine how the number of full-rank epochs affects model accuracy. Table 6 presents the results. SVD-based low-rank training is evaluated on the CIFAR-10 and CIFAR-100 benchmarks. For convenience, we define $\phi=(D-I)/E$ as the ratio of full-rank epochs to the total training budget. When $\phi=0.0$, it corresponds to conventional low-rank training; when $\phi=1.0$, it becomes full-rank training. On both benchmarks, accuracy starts to drop when $\phi$ goes below 0.5, particularly in the range $\phi\in[0.3,0.5]$. We therefore recommend starting with $\phi=0.5$ and adjusting downward as needed.
@@
Table 6: Ablation study on how $\phi=(D-I)/E$ affects the performance of dynamic-rank training. Low-Rank SVD is used for all experiments. When $\phi=0.0$, it becomes the conventional low-rank training. In contrast, when $\phi=1.0$, it becomes full-rank training. The $I$ and $D$ are set to perform the same number of full-rank epochs in the high-noise and low-noise regimes.
@@ 5.3 Computational Cost Analysis
Thus, if $\rho$ is the same, $T_{DR}$ should be higher than $T_{SVD}$ and lower than $T_{F}$. For example, ResNet20 contains 272,762 trainable parameters and the reparameterized model with SVD ($\rho=0.5$) has 155,170 parameters. When $E=150$, $I=60$, and $D=135$, where $\phi=0.5$, $T_{DR}/T_{F}$ becomes $0.7844$, which is shown in Comp. column in Table 3. Our empirical study demonstrates that, by carefully tuning $I$ and $D$, a relatively small DID-I value provides comparable accuracy to the full-rank training. Small DID-I settings in dynamic-rank training makes its training cost more similar to that of the conventional low-rank training.